two identical incandescent light bulbs, each with resistance R=2Ω, are connected to as source with E= 8V and negligible internal resistance. find the current through each bulb is the bulb are connected in series.


1
Expert's answer
2021-05-04T06:44:40-0400

Bulbs are connected in series.

Resistance of each bulb, "R = 2 \Omega"

Total resistance of the circuit, "R_T = 2+2 = 4 \Omega"

Potential in the circuit, "E = 8V"

Total current in circuit, "I = \frac{E}{R_T} = \frac{8}{4} = 2A"

Since bulbs are in series, so current through each bulbs is same i.e. 2A.